# On the role of the unitary transformations in Bell inequalities

### v1

@@ III.3 The violation of the Bell-CHSH inequality in Quantum Field Theory
We have now all tools to discuss the violation of the Bell-CHSH inequality in Quantum Field Theory. Let us proceed with the choice of the Hermitian bounded operators. For such a purpose, we consider the following expressions
@@
where $(f,f^{\prime})\in{\cal O}_{R}$ and $(jf,jf^{\prime})\in{\cal O}_{L}$ are Alice’s and Bob’s test functions, respectively. It is easily checked that the above operators are Hermitian. By construction, $(P_{A_{f}},P_{A_{f^{\prime}}})$ commute with $(P_{B_{jf}},P_{B_{jf^{\prime}}})$. Moreover, it turns out that these operators are projectors, a feature which follows from the presence of the vacuum projector |$\ket{0}\bra{0}$ |. In fact,
${P_{A_{f}}}{P_{A_{f}}}$ $=e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}$
$=e^{-i\varphi(f)}\ket{0}\bra{0}e^{i\varphi(f)}={P_{A_{f}}}\;.$ (59)
For further use, it is helpful to introduce the correlation function for the projectors of Eq. (58), i.e
@@
$\langle 0|\;P_{A_{f}}P_{B_{jf}}\;|0\rangle=e^{-\frac{1}{2}||f||^{2}}\;e^{-\frac{1}{2}||f+jf||^{2}}\;e^{-\frac{1}{2}||jf||^{2}}=e^{-2\eta^{2}((1+\lambda^{2})+\lambda)}\;,$ (61)
@@
The effect of the unitary transformations is well captured by Figs. (1) and (2), in which the behaviours of the correlators $\bra{0}\mathcal{C}\ket{0}$ and $\langle 0|\;{\hat{\cal C}}\;|0\rangle$ as a function of the parameters $(\lambda,\eta)$ are reported.
In Fig. (1), one sees the violation of the Bell-CHSH correlator $\bra{0}\mathcal{C}\ket{0}$ before the use of the unitary transformations. In Fig. (2), the effect of the unitaries is exhibited. Here, the blue surface corresponds to the maximum value attained by $\bra{0}\mathcal{C}\ket{0}$: $2.35463$. The orange surface yields the behaviour of $\bra{0}\hat{\mathcal{C}}\ket{0}$. One sees that the orange surface overcomes the blue one.
Figure 1: Violation of the Bell-CHSH inequality before the use of unitary transformations. The orange surface represents the behaviour of $\bra{0}\mathcal{C}\ket{0}$ as function of the parameters $(\lambda,\eta)$. The blue surface is the classical bound: 2
Figure 2: Behaviour of the Bell-CHSH inequality after the use of unitaries. The orange surface represents the behaviour of $\bra{0}\hat{\mathcal{C}}\ket{0}$ as function of the parameters $(\lambda,\eta)$. The blue surface is related to the maximum value attained by $\bra{0}\mathcal{C}\ket{0}$: $2.35463$.
IV Conclusions
In this work, the role of the unitary transformations has been highlighted. Through a few examples, it has been shown that, in Quantum Mechanics, the Bell operators can be cast into the form
$A={\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}^{\dagger}_{A}{\cal U}_{A}\,=\mathbf{1}=\,{\cal U}_{A}{\cal U}^{\dagger}_{A}\;,\qquad A_{0}=A_{0}^{\dagger}\;,\qquad A_{0}^{2}=\mathbf{1}\;,$ (74)
where $A_{0}$ stands for a reference operator. All parameters leading to the violation of the Bell-CHSH inequality can be encoded in the unitary transformation.
A similar pattern holds in Quantum Field Theory. Here, as exhibited by Eqs. (69) and (72), the size of the violation is considerably improved by making use of unitary transformations, Eqs. (71).
Let us add the following remarks:
  • Looking at Eqs. (71), one is led to figure out that a sequence of unitary transformations could be constructed and applied to the starting operators $(A_{f},A_{f^{\prime}})$ and $(B_{jf},B_{jf^{\prime}})$, Eqs. (58). At each step, the size of the violation should be improved a little bit. Perhaps, this procedure might give a practical way of achieving a final violation close to Tsirelson’s bound: $2\sqrt{2}$.
  • The construction of the bounded Hermitian operators presented in Eqs. (58) looks rather interesting, being easily generalized to a whole family of operators. Let, in fact, $A(\varphi(f))$ denote a generic, non-Hermitian bounded operator in the smeared field $\varphi(f)$. Then, the quantity
    ${\cal A}(\varphi(f))=A^{\dagger}(\varphi(f))\;|0\rangle\langle 0|\;A(\varphi(f))\;,$ (75)
    is Hermitian and bounded.
  • Interestingly enough, the above construction applies to Abelian gauge theories as well. Let $V_{\mu}$ be an Abelian gauge field, subject to the gauge transformation
    $V_{\mu}\rightarrow V_{\mu}+\partial_{\mu}\alpha\;,$ (76)
    and let $V(f)$ be the corresponding smeared field [26]
    $V(f)=\int d^{4}x\;V_{\mu}(x)f^{\mu}(x)\;,$ (77)
    where $\{f^{\mu}\}$ are smooth test functions with compact support. Therefore, the operator
    ${\cal P}_{V}(f)=e^{-iV(f)}\;|0\rangle\langle 0|\;e^{iV(f)}\;,$ (78)
    is gauge invariant, besides of being bounded and Hermitian. Needless to say, expression (78) might open the road for investigating the violation of the Bell-CHSH inequality in the vacuum state of Abelian gauge theories.
@@ Acknowledgments
The authors would like to thank the Brazilian agencies CNPq, CAPES end FAPERJ for financial support. S.P. Sorella, I. Roditi, and M.S. Guimaraes are CNPq researchers under contracts 301030/2019-7, 311876/2021-8, and 309793/2023-8, respectively. F.M. Guedes acknowledges FAPERJ for financial support under the contract SEI-260003/007871/2024.

### v2

@@ III.3 The violation of the Bell-CHSH inequality in Quantum Field Theory
We have now all tools to discuss the violation of the Bell-CHSH inequality in QFT. Let us proceed with the choice of the Hermitian bounded operators. For such a purpose, we consider the following expressions
@@
where $(f,f^{\prime})\in{\cal O}_{R}$ and $(jf,jf^{\prime})\in{\cal O}_{L}$ are Alice’s and Bob’s test functions, respectively. It is easily checked that the above operators are Hermitian. By construction, $(P_{A_{f}},P_{A_{f^{\prime}}})$ commute with $(P_{B_{jf}},P_{B_{jf^{\prime}}})$. Moreover, it turns out that these operators are projectors, a feature which follows from the presence of the vacuum projector $|\ket{0}\bra{0}$. In fact,
[…]
It is important to observe that
$\langle 0|\;\left[P_{A_{f}},P_{B_{jf}}\right]\;|0\rangle$ $=\langle 0|\;\left[P_{A_{f}},P_{B_{jf^{\prime}}}\right]\;|0\rangle=0\;,$
$\langle 0|\;\left[P_{A_{f^{\prime}}},P_{B_{jf}}\right]\;|0\rangle$ $=\langle 0|\;\left[P_{A_{f^{\prime}}},P_{B_{jf^{\prime}}}\right]\;|0\rangle=0\;,$ (49)
showing that the operators $(P_{A_{f}},P_{A_{f^{\prime}}})$ commute with $(P_{B_{jf}},P_{B_{jf^{\prime}}})$ in the vacuum state. For further use, it is helpful to introduce the correlation function for the projectors of Eq. (47), i.e.,
@@
[…]
@@
[…]
Figure 1: Behaviour of the Bell-CHSH inequality before (left panel) and after (rigth panel) the use of unitaries.
IV The Proca field in ($1+1$)-dimensions
We now turn to the case of the massive spin-1 Proca field in two-dimensional Minkowski spacetime. The field $V_{\mu}$ is described by the Lagrangian density
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}(V)F^{\mu\nu}(V)+\frac{M^{2}}{2}V^{\mu}V_{\mu}\;,$ (64)
where $F_{\mu\nu}(V)=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}\;$ and the parameter $M$ corresponds to the physical mass of the spin-1 vector field. The quantized field admits the plane-wave mode expansion
$V_{\mu}(t,x)=\int\frac{dk}{2\pi}\frac{1}{2\omega_{k}}[\epsilon_{\mu}(k)a(k)e^{-ik_{\mu}x^{\mu}}+\epsilon_{\mu}(k)a^{\dagger}(k)e^{ik_{\mu}x^{\mu}}]\;,$ (65)
with dispersion relation $\omega_{k}=\sqrt{k^{2}+M^{2}}$. The polarization vectors satisfy the standard conditions
$k^{\mu}\,\epsilon_{\mu}(k)=0,\qquad\epsilon^{\mu}(k)\,\epsilon_{\mu}(k)=-1,\qquad\epsilon_{\mu}(k)\,\epsilon_{\nu}(k)=-\left(\eta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{M^{2}}\right)\;,$ (66)
with $\eta_{\mu\nu}=\textrm{diag}(+,-)$. The creation and annihilation operators fulfill the canonical commutation relations
$=(2\pi)2\omega_{k}\,\delta(k-q),$ (67)
$[a_{k},a_{q}]$ $=[a^{\dagger}_{k},a^{\dagger}_{q}]=0\;.$
A distinctive property of the Proca field is its transversality condition $\partial^{\mu}V_{\mu}=0,$ which implies that only transverse test functions contribute when the field is smeared. For a smooth compactly supported function $f_{\mu}({\bf\it x})$, we define
$V(f)=\int d^{2}{\bf\it x}\,V_{\mu}({\bf\it x})f^{\mu}({\bf\it x})=a_{f}+a^{\dagger}_{f},$ (68)
with the inner product
$a_{f},a^{\dagger}_{g}]=\langle f|g\rangle_{V}=-\int\frac{dk}{2\pi}\frac{1}{2\omega_{k}}f^{*\,\mu}(k)g^{\nu}(k)\left(\eta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{M^{2}}\right)\;.$ (69)
The Fourier transform of the test functions is given by
$f^{\mu}(k)=\int d^{2}{\bf\it x}\,e^{-ik_{\mu}x^{\mu}}f^{\mu}({\bf\it x})\;.$ (70)
Equivalently, in configuration space, one finds
$\langle f|g\rangle_{V}=-\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f^{\mu}({\bf\it x})\left(\eta_{\mu\nu}-\frac{\partial^{{\bf\it x}}_{\mu}\partial^{{\bf\it x}}_{\nu}}{M^{2}}\right)\Delta({\bf\it x}-{\bf\it y})g^{\nu}({\bf\it y})\;,$ (71)
where
$\Delta({\bf\it x}-{\bf\it y})=\int\frac{dk}{2\pi}\frac{1}{2\omega_{k}}e^{-ik_{\mu}(x^{\mu}-y^{\mu})}=H({\bf\it x}-{\bf\it y})+\frac{i}{2}\Delta_{\rm PJ}({\bf\it x}-{\bf\it y})\;,$ (72)
with $H({\bf\it x}-{\bf\it y})$ and $\Delta_{\rm PJ}({\bf\it x}-{\bf\it y})$ denoting the Hadamard and Pauli–Jordan distributions, respectively [cf. Eqs. (19)–(21)].
To analyze the consequences of transversality, let an arbitrary test function $h_{\mu}({\bf\it x})$ be decomposed into transverse and longitudinal parts,
$h_{\mu}=h_{\mu}^{T}+h_{\mu}^{L},\qquad h_{\mu}^{T}=\left(\eta_{\mu\nu}-\frac{\partial_{\mu}\partial_{\nu}}{\partial^{2}}\right)h^{\nu},\qquad h_{\mu}^{L}=\frac{\partial_{\mu}\partial_{\nu}}{\partial^{2}}h^{\nu}\;,$ (73)
where $\partial^{2}=\partial_{0}^{2}-\partial_{1}^{2}.$ Then, the longitudinal component does not contribute to the smeared field:
$V(h)=\int d^{2}{\bf\it x}\,V^{\mu}h_{\mu}=\int d^{2}{\bf\it x}\,V^{\mu}h^{T}_{\mu}=V(h^{T})\;,$ (74)
which shows that only transverse test functions are physically relevant. A convenient realization is
$f_{\mu}(t,x)=\frac{1}{M}(\partial_{1}f,\partial_{0}f)\;,$ (75)
with $f$ a smooth scalar test function. This ensures that $\partial^{\mu}f_{\mu}=\dfrac{1}{M}(\partial_{0}\partial_{1}f-\partial_{1}\partial_{0}f)=0.$ Substituting into the inner product yields
$\langle f|g\rangle_{V}$ $=-\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f^{\mu}({\bf\it x})g_{\mu}({\bf\it y})\Delta({\bf\it x}-{\bf\it y})$
$=-\frac{1}{M^{2}}\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,(\partial_{1}^{{\bf\it x}}f\partial_{1}^{{\bf\it y}}g-\partial_{0}^{{\bf\it x}}f\partial_{0}^{{\bf\it y}}g)\Delta({\bf\it x}-{\bf\it y})$
$=-\frac{1}{M^{2}}\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f({\bf\it x})g({\bf\it y})(\partial_{1}^{{\bf\it x}}\partial_{1}^{{\bf\it y}}-\partial_{0}^{{\bf\it x}}\partial_{0}^{{\bf\it y}})\Delta({\bf\it x}-{\bf\it y})$
$=\frac{1}{M^{2}}\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f({\bf\it x})g({\bf\it y})(-\partial_{0}^{2}+\partial_{x}^{2})\Delta({\bf\it x}-{\bf\it y})\;.$ (76)
Using the Klein–Gordon equation $(\partial_{0}^{2}-\partial_{x}^{2}+M^{2})\Delta({\bf\it x}-{\bf\it y})=0$, this reduces precisely to the scalar-field inner product,
$\langle f|g\rangle_{V}$ $=\int d^{2}{\bf\it x}\,d^{2}{\bf\it y}\,f({\bf\it x})\Delta({\bf\it x}-{\bf\it y})g({\bf\it y})=\langle f|g\rangle\;,$ (77)
demonstrating that, in two spacetime dimensions, the massive Proca field is dynamically equivalent to a scalar field. The reduction follows from the fact that a vector field in $(1+1)$-dimensions carries a single physical degree of freedom. Consequently, the analysis of Bell-CHSH violations parallels the scalar case. Introducing the bounded Hermitian projectors
${\cal P}_{A_{f}}$ $=e^{-iV(f)}\;|0\rangle\langle 0|\;e^{iV(f)}\;,$ (78)
${\cal P}_{A_{f^{\prime}}}$ $=e^{-iV(f^{\prime})}\;|0\rangle\langle 0|\;e^{iV(f^{\prime})}\;,$
${\cal P}_{B_{jf}}$ $=e^{iV(jf)}\;|0\rangle\langle 0|\;e^{-iV(jf)}\;,$
${\cal P}_{B_{jf^{\prime}}}$ $=e^{iV(jf^{\prime})}\;|0\rangle\langle 0|\;e^{-iV(jf^{\prime})}\;,$
and applying the Tomita-Takesaki modular construction, we obtain for the Bell-CHSH correlator in the vacuum
[…]
in complete agreement with the scalar-field result [cf. Eq. (59)]. Furthermore, once the additional unitary transformations of Eq. (61) are incorporated, the violation is enhanced to the same value reported in Eq. (62).
Thus, in $(1+1)$-dimensions, the Proca field exhibits no qualitative distinction from the massive scalar field regarding Bell-CHSH violations as expected from their duality relation: both theories possess the same physical content, leading to identical Bell-CHSH violations in the vacuum and exhibiting the same enhancement under unitary deformations.
V Conclusions
In this work, we have investigated the role of unitary transformations in optimizing violations of the Bell-CHSH inequality across both quantum-mechanical and field-theoretic settings. The central result is that unitary transformations provide a systematic means of enlarging the parameter space of Bell observables, thereby enhancing the extent of nonlocal correlations extracted from a given system.
At the quantum-mechanical level, we emphasized that Bell observables can always be generated from a fixed reference operator by unitary conjugation, and that the parameters introduced by these unitaries directly control the violation. This unifies the treatment of several well-known examples, including singlet states, GHZ states, and coherent-state entanglement, under a common framework.
Extending this approach to relativistic QFT, we employed the Tomita–Takesaki modular theory to construct localized, bounded Hermitian operators supported in spacelike-separated wedge regions. For the real scalar field in $(1+1)$ dimensions, we explicitly computed the Bell-CHSH correlator in the vacuum, recovering a violation of magnitude $\langle 0|\mathcal{C}|0\rangle=2.35463$. Introducing unitary deformations raised this value to $\langle 0|\hat{\mathcal{C}}|0\rangle=2.54066$, significantly closer to the Tsirelson bound $2\sqrt{2}$.
A novel element of our analysis was the inclusion of the Proca field in $(1+1)$ dimensions. In this case, the transversality condition restricts the dynamics to a single degree of freedom, and the Proca field admits a dual representation in terms of a massive scalar. Concretely, the Proca field in two dimensions admits a Hodge dual representation in terms of the derivative of a scalar, and conversely the scalar can be reconstructed from the dual field tensor. This establishes an exact equivalence between the free Proca and scalar theories, ensuring that their Hilbert spaces and correlation structures coincide. Our Bell-CHSH analysis makes this correspondence fully explicit: the correlator values obtained for the Proca field reproduce those of the scalar theory, both before and after unitary enhancement. This constitutes, to our knowledge, the first demonstration of Bell-CHSH violations in a vector-field vacuum, and it highlights how field-theoretic dualities manifest themselves in the entanglement structure probed by Bell operators.
Our findings highlight two important messages. First, the entanglement structure of the QFT vacuum is sufficiently robust to ensure Bell-CHSH violations across both scalar and vector fields, in line with their duality relations. Second, unitary transformations act as a powerful optimization tool, systematically enhancing these violations and providing a route toward approaching Tsirelson’s bound in field-theoretic settings.
Several directions for future work naturally follow. It would be of interest to extend the present construction to higher-dimensional spacetimes, to interacting field theories, and to gauge-invariant sectors where additional constraints arise. In particular, Abelian and non-Abelian gauge fields provide fertile ground for testing the universality of our findings. A variational approach based on iterated unitary deformations may also provide a systematic method for identifying optimal Bell observables in general QFTs.
In summary, our results reinforce the deep interplay between algebraic structures, unitarity, and entanglement in quantum field theory. They also suggest that unitary freedom remains an underexplored resource for probing the ultimate limits of quantum nonlocality in relativistic systems.
@@ Acknowledgments
The authors would like to thank the Brazilian agencies CNPq, CAPES end FAPERJ for financial support. S. P. Sorella, I. Roditi, and M. S. Guimaraes are CNPq researchers under contracts 301030/2019-7, 311876/2021-8, and 309793/2023-8, respectively. A. F. Vieira is supported by a postdoctoral grant from FAPERJ in the Pós-doutorado Nota 10 program, grant No. E-26/200.135/2025. F. M. Guedes acknowledges FAPERJ for financial support under the contract SEI-260003/007871/2024. D.O.R. Azevedo acknowledges PEDECIBA for financial support.